OpenStudy (anonymous):

write an equation for the parabola vertex (9,4), focus (9,12)

OpenStudy (amistre64):

a = 8 soo and this opens to the .....left y^2 = 4ax

OpenStudy (amistre64):

opps, opens up lol

OpenStudy (amistre64):

4ay = x^2 4(8)y = x^2 32y = x^2 then account for the offcenter

OpenStudy (amistre64):

32(y-4) = (x-9)^2 y = (1/32)(x-9)^2 +4

OpenStudy (anonymous):

Use \[(x-h)^{2} = 4p(y-k) \] where h is the x coordinate of the focus or vertex (both are the same), k is the y coordinate of the vertex and p is the difference between the two y values so... \[(x-9)^{2} = 4(8)(y-4) \] simplifies to \[x ^{2} -18x +81 = 32y - 128\] clean it up to get \[y = 1/32(x ^{2} -18x + 209)\]